# SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback

### v2

@@ C.3 Improvement Recommendations: Revised Prompt
 Gwen, 8, 70, 15
**Task Requirements:**
Analyze the table to identify the tallest penguin based on the ’Height (cm)’ column.
**Problem Rule Explanation:**
Compare the numerical values in the ’Height (cm)’ column to determine the maximum height.
**Problem Rule Application:**
Identify the maximum value from the ’Height (cm)’ column.
**Result Verification:**
Ensure that the identified height corresponds to the tallest penguin.
Appendix D Examples of Generated Data
@@
Table 11: Examples of Generated Data for BIG-Bench tasks
Tasks Generated Data
Causal Judgment In a small town, there is a bakery that makes the best pastries. Every morning, the bakery opens at 7:00 am, and a line of customers forms outside. The bakery owner has a rule that only one person can enter at a time to maintain order. One day, two customers, Alice and Bob, arrive at the same time. Alice follows the rule and waits outside, but Bob ignores the rule and enters the bakery while another customer is still inside. The bakery becomes overcrowded, and a shelf of pastries falls over, ruining the day’s batch. Did Bob cause the pastries to be ruined?
Geometry Generated question: This SVG path element This SVG path element <path d="M 50.00,30.00 L 66.18,35.09 L 72.45,50.00 L 66.18,64.91 L 50.00,70.00 L 33.82,64.91 L 27.55,50.00 L 33.82,35.09 L 50.00,30.00"/>
Object Counting Generated question: I have two violins, a drum, a piano, a flute, and a trumpet. Additionally, I have a cat, a rabbit, a dog, a chicken, and a goat. How many musical instruments do I have?
Epistemic Premise: Olivia suspects that Ethan recognizes that a group of musicians gather in a park, tuning their instruments as the sun sets behind the city skyline. Hypothesis: Ethan recognizes that a group of musicians gather in a park, tuning their instruments as the sun sets behind the city skyline.
Temporal Today, Alex attended several events. Between what times could he have gone to the gym? We know that: Alex had breakfast at 8am. He attended a meeting from 9am to 11am. He was seen at the art gallery from 11am to 1pm. He had lunch with friends from 1pm to 2pm. He was at the cinema from 2pm to 4pm. He visited his grandmother from 4pm to 6pm. The gym closes at 10pm. Between what times could Alex have gone to the gym?
Penguins Generated question: Here is a table where the first line is a header and each subsequent line is a penguin: name, age, height (cm), weight (kg) Louis, 7, 50, 11 Bernard, 5, 80, 13 Vincent, 9, 60, 11 Gwen, 8, 70, 15 For example: the age of Louis is 7, the weight of Gwen is 15 kg, the height of Bernard is 80 cm. What is the age of Vincent?
Appendix E Detailed proof of Theorem 3.3
@@
Because the surrogate loss upper-bounds the 001111 loss point-wise,
@@
where we abbreviated $J(\psi)\!:=\!\mathbb{E}_{q_{\psi}}L\!\bigl{(}f(p,\tilde{x}),\tilde{y}\bigr{)}$.
2. Reduce to the near-optimal generator.
Let $\psi^{\star}$ be any generator that $\varepsilon$-maximises the regularised objective,
@@
Applying the risk-alignment inequality to $\psi^{\star}$,
@@
which is exactly the bound claimed in Theorem 3.3. $\square$

### v3

@@ Appendix E Examples of Generated Data
Gwen, 8, 70, 15. For example: the age of Louis is 7, the weight of Gwen is 15 kg, the height of Bernard is 80 cm. Which is the youngest penguin??
Answer: Bernard
@@
Table 17: Examples of Generated Causal Judgement Data for BIG-Bench tasks
Difficulty Level/Iteration Generated Data
1 In a small town, there is a bakery known for its delicious pastries. The bakery owner, Mr. Thompson, has a strict policy that only his trained staff can bake the pastries. However, one day, a customer named Sarah, who has some baking experience, decides to sneak into the kitchen while the staff is busy. She bakes a batch of pastries without permission. The next morning, Mr. Thompson discovers that the pastries are not up to his usual standards. Did Sarah cause the pastries to be of poor quality?
2 In a bustling restaurant, the head chef, Chef Maria, has a rule that only she and her sous chefs are allowed to create new dishes. One evening, a waiter named Tom, who has a passion for cooking, decides to experiment and prepares a new dish during a slow shift. The next day, Chef Maria tastes the dish and finds it lacks flavor and presentation. Did Tom cause the dish to be of poor quality
3 In a high-tech company, the lead engineer, Ms. Johnson, has a strict protocol that only certified technicians are allowed to perform software updates on the main server. One afternoon, a junior intern named Alex, eager to impress, decides to bypass the protocol and updates the software without authorization. The next day, the server experiences significant downtime due to compatibility issues. Did Alex cause the server downtime?
4 In a prestigious university, the dean, Dr. Roberts, has a clear policy that only faculty members are allowed to supervise student research projects. However, a graduate student named Emily, frustrated with the slow progress of her project, decides to take matters into her own hands and conducts an experiment without faculty approval. When the results are presented, they are deemed unreliable and flawed. Did Emily cause the research results to be unreliable?
5 In a renowned art gallery, the curator, Ms. Patel, has a strict guideline that only professional artists can submit their work for exhibition. One day, a passionate art student named Jake, feeling confident in his skills, sneaks in his painting and hangs it on the wall without permission. When the exhibition opens, visitors notice that Jake’s painting lacks the finesse and depth of the other artworks. Did Jake cause the negative reception of his painting?
6 In a large corporate office, the HR manager, Ms. Lee, has a strict policy that only trained personnel are allowed to conduct interviews for new hires. One day, a receptionist named Mark, eager to help, decides to conduct an interview for a candidate without permission. The next day, during the hiring meeting, it becomes clear that the candidate is not a good fit for the company culture. Did Mark cause the poor hiring decision?
7 In a cutting-edge research lab, the lead scientist, Dr. Nguyen, has a strict rule that only authorized researchers can handle the sensitive equipment. One afternoon, a visiting scholar named Rachel, excited to contribute, decides to use the equipment without permission. The next day, Dr. Nguyen discovers that the data collected is corrupted and unusable. Did Rachel cause the data corruption?
8 In a competitive publishing house, the editor-in-chief, Ms. Carter, has a strict policy that only experienced editors can approve manuscripts for publication. One day, a new intern named David, eager to make a mark, decides to approve a manuscript without consulting anyone. When the book is published, it is filled with errors and inconsistencies. Did David cause the poor quality of the published book?
9 In a prestigious law firm, the managing partner, Mr. Stevens, has a strict rule that only licensed attorneys can represent clients in court. One day, a paralegal named Lisa, wanting to prove her capabilities, decides to represent a client during a hearing without authorization. The next day, the judge dismisses the case due to Lisa’s lack of legal knowledge and experience. Did Lisa cause the dismissal of the case?
10 In a leading aerospace company, the project director, Mr. Carter, has a strict rule that only certified engineers can work on the aircraft design. However, his enthusiastic neighbor, Jake, who has no formal training, often offers unsolicited advice on the design process. One day, Jake manages to convince Mr. Carter to incorporate a design feature he believes will improve aerodynamics. Unfortunately, the feature is flawed and leads to a critical failure during a test flight, resulting in the loss of the prototype. Did Jake cause the failure of the prototype?
Appendix F Detailed proof of Theorem 3.3
@@
Because the surrogate loss upper‑bounds the 011 loss point‑wise,
@@
[…]
2. Reduce to the near‑optimal generator.
Let $\psi^{\star}$ be any generator that $\varepsilon$‑maximises the regularised objective,
@@
Applying the risk‑alignment inequality to $\psi^{\star}$,
@@
which is exactly the bound claimed in Theorem $3.$ $\square$